OpenStudy (anonymous):

True or false. The y-intercept of the straight line with equation Ax + By + C = 0 is -C/B (B ≠ 0).

OpenStudy (marziajackson$):

I would have to say true

OpenStudy (anonymous):

could you elaborate on your answer, you dont have to answer it but just guide me in the right direction...sorry if i sound too picky, i just want to understand the problem too

OpenStudy (anonymous):

Ax+By+c=0 Ax+By=-c divide by -c \[\frac{ x }{\frac{ -c }{A } }+\frac{ y }{ \frac{ _-C }{B }}=1\]

OpenStudy (anonymous):

x intercept is -C/A y intercept is -C/B

OpenStudy (anonymous):

ohhh i get it now, thank you!!

OpenStudy (marziajackson$):

...........